OpenStudy (anonymous):

1. How much sugar is in a 1 liter bottle of soda that is 15 % sugar ? 2. If you grew from 150 cm to 160 cm during one year, what was your percent increase in height ? 3. I have been trying to lose weight. If my diet resulted in a change from 225 lbs. to 210 lbs., what was my percent loss of weight ? 4. What is the percentage of salt in a solution if you mixed 10 ml of 5% salt solution with 5 ml of water ? 5. What is the percentage of salt in a solution if you mixed 10 ml of 5% salt solution with 5 ml of pure salt ?

OpenStudy (anonymous):

1. 15% of 1 litre = 15/100* 1000mL = 50mL 2. increase = 160 - 150 = 10cm increase 10 cm percent increase = ---------- * 100% = ------- * 100% = 6.67% (approx.) original 150 cm

OpenStudy (anonymous):

3. weight loss = 225 - 210 = 15 lb decrease 15 lb percent decrease = ---------- * 100% = ------- * 100% = 6.67% (approx.) original 225 lb 4. What is the percentage of salt in a solution if you mixed 10 ml of 5% salt solution with 5 ml of water ? 4. 5% salt solution means that in 100ml solution, there is 5 ml salt so in 10 ml solution, there is 0.5 ml salt on adding 5 ml water the total volume becomes 15 ml now there is 0.5 ml salt in 15 ml solution 0.5 so percentage of salt = ----- * 100% = 3.33 % 15